The PTO is selling raffle tickets to raise money for classroom supplies. A raffle ticket costs $4. There is 1 winning ticket out of the 110 tickets sold. The winner gets a prize worth $82. Round your answers to the nearest cent.

What is the expected value (to you) of one raffle ticket? $

Calculate the expected value (to you) if you purchase 10 raffle tickets. $

What is the expected value (to the PTO) of one raffle ticket? $

If the PTO sells all 110 raffle tickets, how much money can they expect to raise for the classroom supplies? $
Answer #1

1)expected value (to you) of one raffle ticket =expected win-expected cost =82*(1/110)-4= $-3.25

2) expected value (to you) if you purchase 10 raffle tickets =82*(10/110)-40= $ -32.55

3) expected value (to the PTO) of one raffle ticket = $ 3.25

4) expected to raise for the classroom supplies =110*4-82 =358